OpenStudy (anonymous):

quick computing company produces calculators they have found that the cost, c(x),of making x calculators is a quadratic function in terms of x. the company also discovered that it costs 45 to produce 2 calculators, 81 to produce 4 and 285 to produce 10,whats it cost to produce 6 calculators

OpenStudy (anonymous):

quadratic functions are y = ax^2 + bx + c so you have 3 equations to solve for the three variables 45 = 4a + 2b + c 81 = 16a + 4b + c 285 = 100a + 10b + c solve the system of equation the variables should be a = 2, b= 6, c=25 so y = 2* 6*6 + 6*6 + 25 therefore the total cost is $133